# ExPath: Targeted Pathway Inference for Biological Knowledge Bases via Graph Learning and Explanation

### v1

@@ 6. Conclusion
We introduced ExPath, a novel framework for understanding targeted pathways within biological database. ExPath integrates (1) a protein language model (pLM) for encoding AA-seqs into graph features, (2) PathMamba, a hybrid model to capture local and global dependencies, and (3) PathExplainer, a subgraph learning module that identifies key nodes and edges via trainable pathway masks. We also introduced ML-oriented biological evaluations and a new metric.The experiments involving 301 bio-networks evaluations demonstrated that pathways inferred by ExPath maintain biological meaningfulness. We will release all code and curated 301 bio-network data to facilitate reproducibility and enable future research in data-specific bio-network inference. Future work will expand ExPath to analyze other types of bio-networks, enabling broader applications in systems biology and medicine.
@@
Table 4. Summary of pathway data across four pathway classes.
@@ Appendix A Graph isomorphism and WL test
Weisfeiler-Lehman (WL) test is a widely used graph isomorphism test that forms the foundation of many GNNs. In the 1-WL framework, each node’s representation is iteratively updated by aggregating information from its neighboring nodes, followed by a hashing process to capture the structural patterns of the graph. GNNs leveraging this concept, such as Graph Convolutional Networks (GCNs) and Graph Attention Networks (GATs), essentially perform a similar neighborhood aggregation, making them as expressive as the 1-WL test in distinguishing non-isomorphic graphs (Xu et al., 2019). Modern GNN architectures adhere to this paradigm, making the 1-WL a standard baseline for GNN expressivity.
A.1. Theoretical Analysis
We analyze our explained pathways based on the Weisfeiler-Lehman (WL) tests and expressiveness.
Lemma 0.
(Expressiveness for explanations). When combined with higher expressive models (e.g., it distinguishes more graphs), PathExplainer can generate more finely differentiated (and potentially more “faithful”) explanation pathways (subgraphs). In contrast, a less expressive models merges different graphs (or nodes/substructures) into larger equivalence classes, leading to non-unique, less granular explanations.
@@ A.1. Theoretical Analysis
Let $f:\mathcal{G}\to\mathbb{R}^{k}$ be a GNN-based model, and let $\sim$ denote the equivalence relation induced by $f$, i.e.,
$G\sim H\quad\Longleftrightarrow\quad f(G)=f(H).$
Consider a GNNExplainer objective given by
$\mathcal{L}(E;f,G)=\alpha\cdot D\bigl{(}f(G),f(G\setminus E)\bigr{)}+\beta\cdot\Omega(E),$
where $E\subseteq\text{components}(G)$ (e.g., a subset of edges/nodes/features), $D(\cdot,\cdot)$ is a distance or divergence measure between outputs, and $\Omega(E)$ is a regularization term encouraging compactness or sparsity. Define the explainer’s solution by
$E^{*}(G)\;=\;\underset{E\subseteq G}{\mathrm{argmin}}\;\mathcal{L}(E;f,G).$
@@
If $G\sim H$, then any pair of optimal explanations $E^{*}(G)$ and $E^{*}(H)$ must yield the same minimum objective value. Consequently, there is no unique explanation across $G$ and $H$ within the same equivalence class.
Since $G\sim H$, we have $f(G)=f(H)$. By definition of $E^{*}(G)$,
$\mathcal{L}\bigl{(}E^{*}(G);f,G\bigr{)}\;=\;\min_{E\subseteq G}\;\mathcal{L}(E;f,G).$
Similarly, for $H$,
$\mathcal{L}\bigl{(}E^{*}(H);f,H\bigr{)}\;=\;\min_{E\subseteq H}\;\mathcal{L}(E;f,H).$
Because $f(G)=f(H)$, the distance term $D\bigl{(}f(G),f(G\setminus E)\bigr{)}$ behaves the same as $D\bigl{(}f(H),f(H\setminus E)\bigr{)}$ for corresponding substructures $E$ (insofar as the GNN does not distinguish between the same substructure in $G$ and that in $H$). Furthermore, if $G$ and $H$ are isomorphic under the viewpoint of $f$, any subgraph $E_{G}$ of $G$ can be mapped to a subgraph $E_{H}$ of $H$ with equivalent contribution to $D(\cdot,\cdot)$ and to $\Omega(\cdot)$.
Hence, there exists a subgraph $E_{H}\subseteq H$ such that
$\mathcal{L}(E_{H};f,H)=\mathcal{L}(E^{*}(G);f,G).$
Likewise, one can construct a subgraph $E_{G}$ from $E^{*}(H)$ with the same objective value. Thus, for any $G\sim H$,
$\mathcal{L}\bigl{(}E^{*}(G);f,G\bigr{)}=\mathcal{L}\bigl{(}E^{*}(H);f,H\bigr{)},$
@@
which implies the optimal explanations are not uniquely determined beyond the equivalence class $[G]$ (the set of all graphs equivalent to $G$). In other words, the GNNExplainer’s solution within an equivalence class $\sim$ cannot uniquely distinguish between graphs $G$ and $H$ such that $G\sim H$. Less expressive models merges different graphs (or nodes/substructures) into larger equivalence classes, leading to non-unique, less granular explanations. In contrast, PathExplainer can generate more finely differentiated explanation subgraphs, when combined with higher expressive models. ∎
Lemma 0 (Comparison with k-WL test).
Proof.
The proof of this theorem directly comes from the recent work (Tonshoff et al., 2023; Behrouz and Hashemi, 2024). They prove a similar theorem using 1-d CNN (Tonshoff et al., 2023) or SSM (Behrouz and Hashemi, 2024) with random sampled subgraph. ∎
Lemma 0 (Comparison with 1-WL test).
PathMamba is strictly more expressive than 1-WL GNNs.
Proof.
We first note that PathMamba contains the GIN as a sub-module, which has the same expressive power as the 1-WL test (Xu et al., 2019). Therefore, PathMamba is at least as expressive as 1-WL GNNs. By Lemma 2, there are graphs that cannot be distinguished by $1$-WL GNNs, but can be distinguished by PathMamba. Consequently, PathMamba is strictly more expressive than 1-WL GNNs. ∎
Theorem 4 (Explanations of ExPath).
Based on Lemma 1, 2, and 3, ExPath can generate more finely differentiated (and potentially more “faithful”) explanation pathways (subgraphs) than 1-WL GNN-based methods, and not bounded by any WL GNN methods.
Appendix B Computational Complexity Analysis
@@
Given $K$ tokens, the complexity of Mamba (Gu and Dao, 2024) is linear with respect to $K$. For $m\geq 1$, for each node $v\in V$, we generate $|V|$ sampled pathways with length $m$, the time complexity of global module mamba would be:
@@
since we have $O(|V|\times m)$ tokens. Our PathMamba is faster than the quadratic time complexity $O(|V|^{2})$ of graph transformers (Rampášek et al., 2022).
In practice, combined with GNN, which requires $O(|V|+|E|)$ time, the totall complexity would be:
@@ Appendix C Preprocessing
We searched for and downloaded all the raw data for the human pathway network (referred to as Homo Sapiens in most bio-databases) using KEGG APIs. The data underwent a series of preprocessing steps to ensure its quality and relevance. (1) First, we ensured that all protein nodes in the network were linked to well-characterized genetic origins, specifically their reference amino acid sequence data (Kanehisa et al., 2016). Using KEGG’s gene-to-protein mapping, we filtered the dataset to retain only protein nodes with associated genomic annotations (Kanehisa and Sato, 2020). Protein nodes lacking sequencing data or genetic associations in KEGG were excluded to reduce noise caused by incomplete or ambiguous sequence information. For the retained nodes, their amino acid sequences were extracted and utilized as input features, ensuring a biologically meaningful representation for the learning task. (2) Second, we streamlined the network structure by removing redundant or biologically insignificant interactions. Specifically, we eliminated non-functional self-loops (edges connecting a node to itself without annotated biological relevance) and isolated nodes lacking any edges. This process included both nodes that were initially isolated and those rendered isolated following the first step of node filtering. Since these elements no longer contributed to the network’s connectivity or functional variation, their removal reduced unnecessary complexity and ensured the network focused exclusively on meaningful and biologically interpretable interactions (Bornholdt, 2005). (3) Third, we preserved the edge with properties of protein-protein interactions while removing directional information to transform the network into an undirected graph. This conversion enabled the analysis to emphasize undirected, pairwise interactions, which are often more pertinent to network-based studies, such as clustering, community detection, or functional enrichment analyses. (4) Finally, the pathways were organized into functional classes based on their KEGG pathway labels, ensuring that biologically related pathways were grouped together. The original labels for Environmental Information Processing and Genetic Information Processing were combined into a unified class, Molecular and Cellular Processes, to reflect their shared biological roles in cellular signaling, communication, and gene regulation (Aoki and Kanehisa, 2005). As a result, the pathway data used in this study was categorized into four main classes: Human Diseases, Metabolism, Molecular and Cellular Processes, and Organismal Systems.
Appendix D Baseline
@@
We collected baselines from both message-passing GNNs and more advanced graph models. Graph Convolutional Network (GCN) (Kipf and Welling, 2017) serves as a foundational GNN, leveraging spectral graph theory for node feature aggregation. GraphSAGE (Hamilton et al., 2017) improves scalability by introducing neighborhood sampling and learnable aggregation functions. Graph Attention Network (GAT) (Veličković et al., 2018) incorporates attention mechanisms to assign different importance to neighbors during feature aggregation. Graph Isomorphism Network (GIN) (Xu et al., 2019) achieves high expressivity, distinguishing graph structures with a focus on injective neighborhood aggregation. GPS (Rampášek et al., 2022) combines GNNs with transformer-style global attention to effectively process both local and global graph structures. Similarly, Graph-Mamba (Wang et al., 2024) processes local structures using GNNs and leverages the Mamba module to capture global node relationships.
@@
For explainer baselines, among statistical methods, Random Sampling (RRS) (Kashtan et al., 2004b) serves as a simple baseline by selecting nodes or edges randomly for comparison. Personalized PageRank (PPR) (Iván and Grolmusz, 2011) computes node importance by incorporating a teleportation mechanism that biases the random walk towards specific nodes, effectively capturing both local and global graph structures. Minimum Dominating Set (MDS) (Nacher and Akutsu, 2016; Wuchty, 2014) identifies a minimal set of nodes that can collectively influence or dominate all other nodes in the graph, providing insights into critical nodes for coverage or control. Notably these three statistical methods only do not use AA sequence node features. For gradient-based methods, Saliency (Simonyan et al., 2013) highlights features based on the magnitude of input gradients. InputXGradient (Shrikumar et al., 2017) combines input features with their gradients to capture feature significance. Deconvolution (Mahendran and Vedaldi, 2016) focuses on reconstructing important input features, emphasizing positive influences. ShapleyValueSampling (Strumbelj and Kononenko, 2010) estimates feature importance using a game-theoretic approach. GuidedBackpropagation (Springenberg et al., 2014) refines gradients to highlight only relevant activations. From GNN-specific explainability approaches, we adopted GNNExplainer (Ying et al., 2019) to uncover subgraphs and features that are critical for predictions, and PGExplainer (Luo et al., 2020), which uses a neural network to identify significant graph components.
Appendix E Metrics
@@
As a classic method of biological functional enrichment analysis, GO analysis evaluates whether specific biological processes, molecular functions, or cellular components are statistically overrepresented in a given set of genes (i.e., gene nodes from subgraphs) compared to a background gene set (Huang et al., 2009). The results provide a list of GO terms that highlight the biological functions most significantly represented in the input gene nodes (Ashburner et al., 2000). In our study, we conducted GO enrichment analysis using the R package clusterProfiler (Yu et al., 2012) to identify enriched GO terms.
To assess Breadth of the subgraphs’ biological functions, we use the Number of Enriched Biological Functions (#EBF) as a metric. For an input subgraph $S_{i}$, #EBF is defined as:
#$\#EBF(S_{i})=|GO_{\text{enriched}}(S_{i})|,$
@@
To assess Depth of the subgraphs’ biological functions, we used the Enrichment Contribution Score (ECS) as a metric. The ECS evaluates the relative contribution of the top-weighted genes, denoted as $G_{\text{Top}}$, to the enrichment of biological functions. The ECS can be assessed by following steps: Let $G=\{g_{1},g_{2},\dots,g_{n}\}$ represent the ranked list of gene nodes, sorted by their importance scores (weights) $w=\{w_{1},w_{2},\dots,w_{n}\}$, where $w_{1}\geq w_{2}\geq\dots\geq w_{n}$. Define $G_{\text{Top}}=\{g_{1},g_{2},\dots,g_{\text{Top}}\}$ to include only genes with the top weights, selected based on a ratio $R\%$ (defaulted as 30%), as a subset of $G$. Then, perform GO analysis based on $G_{\text{Top}}$ for each input subgraph $S_{i}$. The ECS is calculated as the average number of enriched items for each gene in $G_{\text{Top}}$ across all subgraphs $S_{i}$, and is defined as:
@@
To assess Reliability of the subgraphs’ biological functions, we use the well-established statistical concept P-value as a metric. Specifically, the P-value is calculated as the average of the P-values from statistical tests performed for biological function enrichment in each subgraph above. Since during the GO analysis, we accept the item only with a P-value lower than 0.05, the average P-value reported here is naturally lower than this threshold. A lower average P-value indicates greater reliability in the enrichment results across the subgraphs.
Appendix F Ablation Study
F.1. Classifier ablation
Table 5. PathExplainer fidelity score comparison with other classifier models.

### v2

@@ Conclusion
We introduced ExPath, a novel framework for understanding targeted pathways within biological knowledge bases. ExPath integrates PathMamba, a hybrid model to capture local and global dependencies; and PathExplainer, a subgraph learning module that identifies key nodes and edges via trainable pathway masks. ExPath seamlessly integrated biological foundation models to encode the experimental molecular data. We also introduced machine-learning-oriented biological evaluations and a new metric. The experiments involving 301 bio-networks evaluations demonstrated that pathways inferred by ExPath maintain biological meaningfulness. Future work will expand ExPath to analyze other types of bio-networks, enabling broader applications in systems biology and medicine.
@@
Table 5: Summary of pathway data across four pathway classes.
@@ Appendix A Graph isomorphism and WL test
Weisfeiler-Lehman (WL) test is a widely used graph isomorphism test that forms the foundation of many GNNs. In the 1-WL framework, each node’s representation is iteratively updated by aggregating information from its neighboring nodes, followed by a hashing process to capture the structural patterns of the graph. GNNs leveraging this concept, such as Graph Convolutional Networks (GCNs) and Graph Attention Networks (GATs), essentially perform a similar neighborhood aggregation, making them as expressive as the 1-WL test in distinguishing non-isomorphic graphs (Xu et al. 2019). Modern GNN architectures adhere to this paradigm, making the 1-WL a standard baseline for GNN expressivity.
Appendix B Proof
[…]
@@ Appendix B Proof
Given equivalent graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, let $f:\mathcal{G}\to\mathbb{R}^{k}$ be a GNN-based model, and let $\sim$ denote the equivalence relation induced by $f$, i.e.,
$\mathcal{G}_{1}\sim\mathcal{G}_{2}\quad\Longleftrightarrow\quad f(\mathcal{G}_{1})=f(\mathcal{G}_{2}).$
We define an explanatory subgraph $\hat{\mathcal{G}}\subseteq\mathcal{G}$ and consider a PathMamba objective given by
$\mathcal{L}\bigl{(}\hat{\mathcal{G}};f,\mathcal{G}\bigr{)}\;=\;\alpha\,D\bigl{(}f(\mathcal{G}),f(\hat{\mathcal{G}})\bigr{)}\;+\;\beta\,\Omega\bigl{(}\hat{\mathcal{G}}\bigr{)}$,
where $D(\cdot,\cdot)$ is a divergence between model outputs and $\Omega(\cdot)$ penalises the size or complexity of $\hat{\mathcal{G}}$. The optimal explanation for $\mathcal{G}$ is
$E(\mathcal{G})\;=\;\operatorname*{arg\,min}_{\hat{\mathcal{G}}\subseteq\mathcal{G}}\mathcal{L}\!\bigl{(}\hat{\mathcal{G}};f,\mathcal{G}\bigr{)}.$
@@
If $\mathcal{G}_{1}\sim\mathcal{G}_{2}$, then any pair of optimal explanations $E(\mathcal{G}_{1})$ and $E(\mathcal{G}_{2})$ must yield the same minimum objective value. Consequently, there is no unique explanation across $\mathcal{G}_{1}$ and $H$ within the same equivalence class. Since $\mathcal{G}_{1}\sim\mathcal{G}_{2}$, we have $f(\mathcal{G}_{1})=f(\mathcal{G}_{2})$. By definition of $E(\mathcal{G}_{1})$,
$\mathcal{L}\bigl{(}E(\mathcal{G}_{1});f,\mathcal{G}_{1}\bigr{)}\;=\;\min_{\hat{\mathcal{G}}_{1}\subseteq\mathcal{G}_{1}}\;\mathcal{L}(\hat{\mathcal{G}}_{1};f,\mathcal{G}_{1}).$
Similarly, for $\mathcal{G}_{2}$,
$\mathcal{L}\bigl{(}E(\mathcal{G}_{2});f,\mathcal{G}_{2}\bigr{)}\;=\;\min_{\hat{\mathcal{G}}_{2}\subseteq\mathcal{G}_{2}}\;\mathcal{L}(\hat{\mathcal{G}}_{2};f,\mathcal{G}_{2}).$
[…]
$\mathcal{L}\bigl{(}E(\mathcal{G}_{1});f,\mathcal{G}_{1}\bigr{)}=\mathcal{L}\bigl{(}E(\mathcal{G}_{2});f,\mathcal{G}_{2}\bigr{)},$
@@
which implies the optimal explanations are not uniquely determined beyond the equivalence class $\mathcal{G}_{1}]$ (the set of all graphs equivalent to $\mathcal{G}_{1}$). In other words, PathExplainer cannot uniquely distinguish between subgraphs $\hat{\mathcal{G}_{1}}$ and $\hat{\mathcal{G}_{2}}$ such that $\mathcal{G}_{1}\sim\mathcal{G}_{2}$. Less expressive models merge different graphs into larger equivalence classes, leading to non-unique, less granular explanations. In contrast, when combined with higher expressive models (i.e. PathMamba), PathExplainer can generate more finely differentiated explanation subgraphs. ∎
Appendix C Computational Complexity Analysis
@@
Given $K$ tokens, the complexity of Mamba (Gu and Dao 2024) is linear with respect to $K$. For $m\geq 1$, for each node $v\in V$, we generate $|V|$ sampled pathways with length $m$, the time complexity of global module mamba would be:
@@
since we have $O(|V|\times m)$ tokens. Our PathMamba is faster than the quadratic time complexity $O(|V|^{2})$ of graph transformers (Rampášek et al. 2022).
In practice, combined with GNN, which requires $O(|V|+|E|)$ time, the total complexity would be:
@@ Appendix D Preprocessing
Data Acquisition.
We searched for and downloaded all the raw data for the human pathway network (referred to as Homo Sapiens in most bio-databases) using KEGG APIs. The data underwent a series of preprocessing steps to ensure its quality and relevance.
(1) Node Filtering and Feature Construction.
We ensured that all protein nodes in the network were linked to well-characterized genetic origins, specifically their reference amino acid sequence data (Kanehisa et al. 2016). Using KEGG’s gene-to-protein mapping, we filtered the dataset to retain only protein nodes with associated genomic annotations (Kanehisa and Sato 2020). Protein nodes lacking sequencing data or genetic associations in KEGG were excluded to reduce noise caused by incomplete or ambiguous sequence information. For the retained nodes, their amino acid sequences were extracted and utilized as input features, ensuring a biologically meaningful representation for the learning task.
(2) Edge and Structure Cleaning.
We streamlined the network structure by removing redundant or biologically insignificant interactions. Specifically, we eliminated non-functional self-loops (edges connecting a node to itself without annotated biological relevance) and isolated nodes lacking any edges. This process included both nodes that were initially isolated and those rendered isolated following the first step of node filtering. Since these elements no longer contributed to the network’s connectivity or functional variation, their removal reduced unnecessary complexity and ensured the network focused exclusively on meaningful and biologically interpretable interactions (Bornholdt 2005).
(3) Graph Conversion
We preserved the edge with properties of protein-protein interactions while removing directional information to transform the network into an undirected graph. This conversion enabled the analysis to emphasize undirected, pairwise interactions, which are often more pertinent to network-based studies, such as clustering, community detection, or functional enrichment analyses.
(4) Functional Categorization.
The pathways were organized into functional classes based on their KEGG pathway labels, ensuring that biologically related pathways were grouped together. The original labels for Environmental Information Processing and Genetic Information Processing were combined into a unified class, Molecular and Cellular Processes, to reflect their shared biological roles in cellular signaling, communication, and gene regulation (Aoki and Kanehisa 2005). As a result, the pathway data used in this study was categorized into four main classes: Human Diseases, Metabolism, Molecular and Cellular Processes, and Organismal Systems.
Appendix E Baselines
E.1 Classifier models.
We collected baselines from both message-passing GNNs and more advanced hybrid graph models.
@@ Appendix E Baselines
  • Message-passing GNNs: Graph Convolutional Network (GCN) (Kipf and Welling 2017) serves as a foundational GNN, leveraging spectral graph theory for node feature aggregation. GraphSAGE (Hamilton, Ying, and Leskovec 2017) improves scalability by introducing neighborhood sampling and learnable aggregation functions. Graph Attention Network (GAT) (Veličković et al. 2018) incorporates attention mechanisms to assign different importance to neighbors during feature aggregation. Graph Isomorphism Network (GIN) (Xu et al. 2019) achieves high expressivity, distinguishing graph structures with a focus on injective neighborhood aggregation.
  • Hybrid graph models: GPS (Rampášek et al. 2022) combines GNNs with transformer-style global attention to effectively process both local and global graph structures. Similarly, Graph-Mamba (Wang et al. 2024) processes local structures using GNNs and leverages the Mamba module to capture global node relationships.
E.2 Explainer methods.
We collected baselines from statistical methods, gradient-based methods, and GNN-specific explainer methods.
@@
  • Statistical methods: Random Sampling (RRS) (Kashtan et al. 2004b) serves as a simple baseline by selecting nodes or edges randomly for comparison. Personalized PageRank (PPR) (Iván and Grolmusz 2011) computes node importance by incorporating a teleportation mechanism that biases the random walk towards specific nodes, effectively capturing both local and global graph structures. Minimum Dominating Set (MDS) (Nacher and Akutsu 2016; Wuchty 2014) identifies a minimal set of nodes that can collectively influence or dominate all other nodes in the graph, providing insights into critical nodes for coverage or control. Notably these three statistical methods only do not use AA sequence node features.
  • Gradient-based methods: Saliency (Simonyan, Vedaldi, and Zisserman 2013) highlights features based on the magnitude of input gradients. InputXGradient (Shrikumar, Greenside, and Kundaje 2017) combines input features with their gradients to capture feature significance. Deconvolution (Mahendran and Vedaldi 2016) focuses on reconstructing important input features, emphasizing positive influences. ShapleyValueSampling (Strumbelj and Kononenko 2010) estimates feature importance using a game-theoretic approach. GuidedBackpropagation (Springenberg et al. 2014) refines gradients to highlight only relevant activations.
  • GNN-specific explainability approaches, we adopted GNNExplainer (Ying et al. 2019) to uncover subgraphs and features that are critical for predictions, and PGExplainer (Luo et al. 2020), which uses a neural network to identify significant graph components.
Appendix F Metrics
@@
  • Diameter: Diameter of a (connected, unweighted) graph $G=(V,E)$ is
    $\operatorname{diam}(G)=\max_{u,v\in V}\;d_{G}(u,v),$
    where $d_{G}(u,v)$ denotes the length (number of edges) of the shortest path between vertices $u$ and $v$. Intuitively, it is the length of the longest among all shortest paths, capturing the farthest distance that must be traversed within the graph. Because each method yields many subgraphs, we report the average diameter:
    $\overline{D}=\frac{1}{|\mathcal{S}|}\sum_{G_{S}\in\mathcal{S}}\operatorname{diam}(G_{S}),$
    i.e. the mean diameter over the set $\mathcal{S}$ of all extracted subgraphs. A larger $\overline{D}$ indicates that the explanations retain longer end-to-end interactions, aligning with pathway-level preservation.
As a classic method of biological functional enrichment analysis, GO analysis evaluates whether specific biological processes, molecular functions, or cellular components are statistically overrepresented in a given set of genes (i.e., gene nodes from subgraphs) compared to a background gene set (Huang, Sherman, and Lempicki 2009). The results provide a list of GO terms that highlight the biological functions most significantly represented in the input gene nodes (Ashburner et al. 2000). In our study, we conducted GO enrichment analysis using the R package clusterProfiler (Yu et al. 2012) to identify enriched GO terms.
  • #EBF: To assess Breadth of the subgraphs’ biological functions, we use the Number of Enriched Biological Functions (#EBF) as a metric. For an input subgraph $S_{i}$, #EBF is defined as:
    $EBF(S_{i})=|GO_{\text{enriched}}(S_{i})|,$
@@
  • ECS: To assess Depth of the subgraphs’ biological functions, we used the Enrichment Contribution Score (ECS) as a metric. The ECS evaluates the relative contribution of the top-weighted genes, denoted as $G_{\text{Top}}$, to the enrichment of biological functions. The ECS can be assessed by following steps: Let $G=\{g_{1},g_{2},\dots,g_{n}\}$ represent the ranked list of gene nodes, sorted by their importance scores (weights) $w=\{w_{1},w_{2},\dots,w_{n}\}$, where $w_{1}\geq w_{2}\geq\dots\geq w_{n}$. Define $G_{\text{Top}}=\{g_{1},g_{2},\dots,g_{\text{Top}}\}$ to include only genes with the top weights, selected based on a ratio $R\%$ (defaulted as 30%), as a subset of $G$. Then, perform GO analysis based on $G_{\text{Top}}$ for each input subgraph $S_{i}$. The ECS is calculated as the average number of enriched items for each gene in $G_{\text{Top}}$ across all subgraphs $S_{i}$, and is defined as:
@@
  • P-value: To assess Reliability of the subgraphs’ biological functions, we use the well-established statistical concept P-value as a metric. Specifically, the P-value is calculated as the average of the P-values from statistical tests performed for biological function enrichment in each subgraph above. Since during the GO analysis, we accept the item only with a P-value lower than 0.05, the average P-value reported here is naturally lower than this threshold. A lower average P-value indicates greater reliability in the enrichment results across the subgraphs.
Appendix G Ablation Study
Table 6: Sequential Mamba module evaluation with LSTM.
Models Overall Accuracy
PathMamba 0.744 ± 0.015
w/ LSTM 0.730 ± 0.014
Figure 6: Comparison of subgraphs extracted from the TCR signaling pathway using two different methods. The TCR Subgraph on the left is from the RSS method, and the TCR Subgraph on the right is from the proposed method. The subgraph nodes and their signaling modules are colored in red. The disruptions within signaling paths are marked in green boxes.
G.1 Sequential module evaluation
Table 6 shows the results when the Mamba module of PathMamba was replaced with an RNN-based model (LSTM). The classification accuracy was 0.730, and PathMamba (0.744) exceeds the LSTM-based model. This suggests that Mamba’s selective state-space modeling is particularly important for identifying complex biological relationships in pathways where both local interactions and long-range functional dependencies influence biological outcomes.
G.2 Classifier evaluation
Table 7: PathExplainer fidelity score comparison with other classifier models.